OpenStudy (anonymous):

The circuit below contains two dependent sources: a voltage controlled voltage source and a voltage controlled current source.

OpenStudy (nurali):

Formula Vout = Vin*(R3-B*R1*R3)/(R3-(B*R1*R3)+R1+R2+A*R2) i1=V0/(R3-B*R1*R3) and Vth is same as Vout. Isc=vIN*(1-B*R1)/(A*R2+R1+R2) Rth=VO/iSC where VO is Vth

OpenStudy (nurali):

If correct please click best response

OpenStudy (anonymous):

If you can see the link http://forum.allaboutcircuits.com/showthread.php?t=46894 for me, Why does he remove the dependent source to solve.